OpenStudy (anonymous):

Find the period of the function. y = 5 cos 1/2 x

OpenStudy (anonymous):

@Butterfly16

OpenStudy (anonymous):

answer choices are 4 pi and 5

OpenStudy (anonymous):

@Yahoo! @zepdrix

OpenStudy (anonymous):

Okay, I'll show you. :o) Hold on

OpenStudy (anonymous):

f(x)= a sin(bx-x)+d to find the period: 2pi/b so 2pi/1/2. Do you know how to solve it? :o)

OpenStudy (anonymous):

not exactly

OpenStudy (anonymous):

When you have two fractions, like this, you take the inverse of 1/2, which is 2/1 or 2 and multiply it by 2pi, which gives you what? :o) \[2*2\pi\]

OpenStudy (anonymous):

so its 4pi?

OpenStudy (anonymous):

\[\frac{ 2\pi }{ \frac{ 1 }{ 2 } } \rightarrow 2\pi *\frac{ 2 }{ 1 }\rightarrow 2\pi*2\] Awesome job! :o)

OpenStudy (anonymous):

thankyou andcan you help me with this last one too? Identify intervals on which the function is increasing, decreasing, or constant. g(x) = 2 - (x - 7)^2

OpenStudy (anonymous):

the answer choices are Increasing: x < -7; decreasing: x > -7 Increasing: x < 7; decreasing: x > 7 Increasing: x > 2; decreasing: x < 2

OpenStudy (anonymous):

I'm not sure how to solve it, but if you graph it, see where the value of x increases and where the value of x starts to decrease.

OpenStudy (anonymous):

i got b

OpenStudy (anonymous):

Well, if you look at it, when the graph is increasing (or going up) x equals these numbers: -7, -2, 1, 2. When the graph hits 2, the x values turn to: 2, 1, -2, -7. See how the value of x increases on the left side of the graph and then starts decreasing again on the right after it turns at 2?